OpenStudy (whalenjacob):

Two lines, A and B, are represented by the following equations: Line A: y = x − 1 Line B: y = −3x + 11 Which of the following options shows the solution to the system of equations and explains why?

OpenStudy (whalenjacob):

i just a answer

OpenStudy (whalenjacob):

@Slenderman

OpenStudy (slenderman):

what are your options

OpenStudy (whalenjacob):

(3, 2), because the point does not lie on any axis (3, 2), because one of the lines passes through this point (3, 2), because the point lies between the two axes (3, 2), because both lines pass through this point

OpenStudy (luffingsails):

For it to be a solution at all... the given point must satisfy both equations.

OpenStudy (slenderman):

im going to say d

OpenStudy (whalenjacob):

ok thanks

OpenStudy (whalenjacob):

i have a couple more if that ok?

OpenStudy (whalenjacob):

Line M is represented by the following equation: x − y = 8 Which equation completes the system that is satisfied by the solution (18, 10)?

OpenStudy (whalenjacob):

2x − y = 26 x + y = 18 2x − 2y = 36 x − y = −28

OpenStudy (luffingsails):

Just take the point (18,10) and put it into each equation. Only one of them will make a true statement (more than likely). Example: x + y = 18 18 + 10 = 18 28 = 18... false. So this one can be thrown out.

OpenStudy (whalenjacob):

so b

OpenStudy (whalenjacob):

Two lines, A and B, are represented by the following equations: Line A: 3x + 3y = 12 Line B: x + y = 4 Which statement is true about the solution to the set of equations?

OpenStudy (whalenjacob):

It is (12, 4). There are infinitely many solutions. It is (4, 12). There is no solution.

OpenStudy (luffingsails):

No, not b at all. I just proved that b could NOT be the answer.

OpenStudy (whalenjacob):

ik im jk

OpenStudy (whalenjacob):

c

OpenStudy (whalenjacob):

@luffingsails